OpenStudy (thomas5267):

Proof the following: \[ \sum_{r=1}^{n}\sin\left(\frac{2\pi r}{n}\right)=0 \\ \sum_{r=1}^{n}\cos\left(\frac{2\pi r}{n}\right)=0 \]

OpenStudy (thomas5267):

Try not use complex numbers if possible.

OpenStudy (sidsiddhartha):

do u know integration?

OpenStudy (sidsiddhartha):

\[u ~can ~substitute ~(r/n)~as ~x\\then ~\it~will ~be\\ \int\limits_{0}^{1}\sin(2 \pi)x~dx\] now just simple integration :)

OpenStudy (thomas5267):

I know what integration is but I don't know how you transformed a finite sum to a integral.

OpenStudy (thomas5267):

The even case is easy to prove but the odd case is hard. @ganeshie8 @ikram002p

OpenStudy (ikram002p):

it will be something like this :- if you start from - then you well end up with + to get a sum of zero , i have to go :'( cant continue